OpenStudy (anonymous):

A circle has area 72π yd². What is the area of a 10° sector of the circle? a. 2π yd² b. 3π yd² c. 4π yd² d. 5π yd² Thanks! x

OpenStudy (anonymous):

\[10^{0}\] is \[\frac{ 1 }{ 36 }\] of 360

OpenStudy (anonymous):

so the area of it will be 1/36 of the whole area

OpenStudy (anonymous):

\[72 \pi \times \frac{ 1 }{ 36 }=2 \pi\]

OpenStudy (anonymous):

I have 22.6??

OpenStudy (anonymous):

I'm gonna try that again

OpenStudy (anonymous):

I got a. 2π

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

Thank you so much!